The table below contains data on individual’s usage of a social media platform. It categorizes the data on the age range of the individual and their reported usage level. Use this table to answer the question that follows it.


Low

Moderate

High

18-30

286

592

464

31-55

209

438

381

56+

254

237

139

What is the probability that a randomly selected individual reports a high level of usage?
Answer #1

Given,

Low Moderate High Total
18 - 30 286 592 464 1342
31-35 209 438 381 1028
56+ 254 237 139 630
Total 749 1267 984 3000
Low Moderate High
18 - 30 286/3000 = 0.0953 592/3000 = 0.1973 464/3000 = 0.1546
31-35 209/3000 = 0.0696 438/3000 = 0.146 381/3000 = 0.127
56+ 254/3000 = 0.0846 237/3000 = 0.079 139/3000 = 0.0463

The probability that a randomly selected individual reports a high level of usage = 0.154+0.127+0.0463

= 0.3273

The probability that a randomly selected individual reports a high level of usage is 0.3273.